OpenStudy (anonymous):

HELP / / / A function is shown below: f(x) = x3 + 4x2 - x - 4 Part A: What are the factors of f(x)? Show your work. (3 points) Part B: What are the zeros of f(x)? Show your work. (2 points) Part C: What are the steps you would follow to graph f(x)? Describe the end behavior of the graph of f(x). (5 points)

OpenStudy (anonymous):

x3 + 4x2 - x - 4 an expression of 4 terms means we want to find either - a perfect square trinomial - two binomials can you see if there is a perfect square trinomial?

OpenStudy (anonymous):

There isn't one? (:

OpenStudy (anonymous):

here is hint: what about.. 4x2 - x - 4 if 4 = 2^2 then 4x2 - x - 4 = (2x)^2 - x - (2)^2